OpenStudy (anonymous):

Word Problem: Jo invested part of her $8000 at 12% and part at 8%. If the annual investment income is $825, how much was invested at each rate?

OpenStudy (anonymous):

total amount=$8000 total Income=$825 let the amount invested at rate 12%be x then amount invested at 8% rate=8000-x time period=1year 825=x*12/100+(8000-x)8/100 12x+64000-8x=82500 4x=18500 x=4625 let the amount invested at rate 12%=$4625 amount invested at 8% rate=8000-x=8000-4625=$3375